OpenStudy (anonymous):

Please help If f(x) = 6x - 2 and g(x)= -3x + 7, then for what value of x is f(x) = g(x)?

OpenStudy (anonymous):

6x-2=-3x+7 9x=9 x=1 at f(x) and g(x)

OpenStudy (petiteme):

f(x) = 6x - 2 and g(x)= -3x + 7 replace the f(x) to g(x) -3x+7=6x-2 transpose 6 to other side and also 7 to the other side -6x - 3x = -7-2 -9x=-9 divide 9 to both sides, so x=-9/-9 x=1 :)

OpenStudy (yttrium):

Why is it under calculus 1???

OpenStudy (anonymous):

YES X=1.